# Techniques for mass peak reconstruction in Searches for Long-Lived Heavy Neutral Leptons Decaying to a Lepton and a $\rho$ Meson

### v1

@@ Abstract
Keywords:
Heavy Neutral Leptons (HNLs) Semileptonic Decay SHiP LHC mass reconstruction
@@ 1 Introduction
These displaced decays provide a unique experimental insight to search for and characterize HNLs. By reconstructing the kinematics of the decay products, searches can directly constrain the HNL mass and mixing parameters, enabling tests of both the seesaw mechanism and models of sterile neutrinos across a broad range of parameter space. Several experimental programs have placed significant constraints on the HNL parameter space across a wide mass range. In the multi-GeV regime, dedicated searches by ATLAS and CMS provide the most stringest limits on active–sterile mixing ATLAS2023 ; CMS2022 .
In the sub-GeV mass range, experiments such as NA62 have recently searched for the decay $K\rightarrow lN$ and set upper limits on its branching fraction, thereby directly constraining HNL production in kaon decays NA62HNL2021 . Additionally, fixed-target experiments like PS191 and CHARM have been reinterpreted in modern global analyses Drewes2018 to constrain active-sterile mixing and mass range up to 2 GeV. The future experiment at a dedicated beam-dump facility at CERN, SHiP, is expected to significantly extend the search sensitivity in the mass range 0.5-5 GeV to smaller mixing angles SHIP2016 . In experimental searches, the reconstructed or inferred HNL mass is used as the final discriminant to enhance signal sensitivity and suppress background contamination. An accurate determination of the HNL mass plays a central role in setting exclusion limits and defining the sensitivity reach of each analysis across different masses and lifetimes.
@@ 5 Performance Evaluation of the $\rho$-Mass Constraint Method in Beam-Dump and Collider Scenarios
To assess the $\rho$-mass constraint method introduced in Section 4, we evaluate its performance across two representative experimental environments: (i) a fixed-target beam-dump scenario, exemplified by the SHiP experiment, and (ii) a high-energy proton-proton collider scenario, such as the LHC. These two classes of experiments differ significantly in terms of HNL production modes, event topologies, and detector geometries. The objective is twofold: first, to verify whether the kinematic constraint method is able to reconstruct a mass peak as closely as possible to the truth HNL mass and also of smaller width than the simple invariant mass $m_{l\pi}$; and second, to test its discriminatory power in differentiating these from $N\to\ell\pi$ that yields the same charged particles final state but lacks the additional $\pi^{0}$. This section presents simulation-based studies for each experimental scenario at particle level, hence without the effects from detector response.
@@ 5.1 Fixed-Target Scenario
Figure 2: Longitudinal neutral–pion momentum $p_{\pi^{0}}^{\parallel}$. The filled (red) histogram shows the nominal (MC truth), while the hatched (blue/green) histograms show the two kinematic solutions $q_{\min}^{\parallel}$ and $q_{\max}^{\parallel}$ obtained with the $\rho$-mass constraint method. The sample corresponds to $N\!\to\!\mu\rho$ with $m_{N,GeN}=1~\mathrm{GeV}$ and $c\tau=50~\mathrm{m}$, generated with EventCalc--SHiP in a SHiP-like configuration. The $q_{\min}^{\parallel}$ solution is softer and peaks at low momentum, whereas $q_{\max}^{\parallel}$ is broader and extends to higher momenta.
@@ 5.2 Collider Scenario
Figure 4: Longitudinal neutral pion momentum $p_{\pi^{0}}^{\parallel}$ for an LHC-like sample with $m_{N,Gen}=2~\mathrm{GeV}$, produced via $pp\!\to\!W^{\pm}\!+\!X\to\ell N\!+\!X$ and $N\!\to\!\ell\rho$. The filled histogram shows the nominal (MC truth), while the hatched histograms correspond to the two kinematic solutions $q_{min/max}^{\parallel}$ obtained with the $\rho$-mass constraint method. The $q_{\min}^{\parallel}$ solution is softer and peaks at low momentum, whereas $q_{\max}^{\parallel}$ is broader and extends to higher momenta.
@@
Figure 6 compares the $m_{N}^{\rho}$ distributions for the $\mu\rho$ and $\mu\pi$ HNL deccay channels. The reconstructed $m_{N}^{\rho}$ forms a displaced, narrower peak for $\mu\pi$ HNL deccay, enabling clear separation of the decay hypotheses. The observed peak separation reflects the method’s sensitivity to the underlying decay structure. At detector level, however, the achievable discrimination will be limited by the track–momentum resolution at the displaced vertex. For example, ATLAS reports percent-level dimuon mass resolution in the $\mathrm{GeV}$ range for prompt, combined muons ATLASMuonCalib2023 ; resolutions for displaced tracks are typically worse and analysis-dependent. Hence, while the particle-level peak offset is plausibly resolvable, a definitive assessment requires full detector simulation. In addition to the $m_{N}^{\rho}$ peak position, there are two other handles to distinguish the $\ell\rho$ and $\ell\pi$ HNL decay channels (in tracking-only reconstruction):
@@ 6 $W$-Mass Constraint Method
As it was mentioned in Section 3 to further reconstruct the kinematics of the HNL decay, it is useful to exploit momentum conservation in an appropriately chosen reference frame, where we follow essentially the method used for $pp\to\ell W,W\to\ell N,N\to\ell\ell\nu$ described in ATLAS2023 , by replacing the $\nu$ by the $\pi^{0}$ in the $N\to\ell\rho,\rho\to\pi\pi^{0}$. In the rotated frame the energy-momentum conservation relation can be written as:
@@
Analogous to the $\rho$-mass constraint, the $W$-mass constraint yields two kinematic solutions for the HNL mass, denoted as $m_{N}^{\text{W,min/max}}$ . The performance of both solutions is studied using MC-simulated events in Sec. 6.2. In some events, however, no real solution exists because the radicand in Equation 9 becomes negative. To address this, we use an adaptive $W$-mass treatment described in Section 6.1, in analogy to Ref. ATLAS2023 for $pp\to\ell W,W\to\ell N,N\to\ell\ell\nu$.
@@ 7 Outlook: Mass Reconstruction of HNLs Coupling to $\tau$ Leptons
Probing HNL couplings to $\tau$ leptons in HNL decays to final states with a tau lepton plus additional particles is compelling yet challenging. Unlike the electron or muon cases, the $\tau$ must be reconstructed through its decays, which reduces sensitivity due to the $\tau$ branching fractions to the used channels. Moreover, for almost all subsequent $\tau$-lepton decays, leptonic or semileptonic, the reconstruction of a mass peak is not possible due to the missing momentum from the emitted neutrino(s). A notable extension of the method applies for $m_{N}>m_{\tau}+m_{\pi}$, namely the channel $N\to\tau^{+}\pi^{-}$ with $\tau^{+}\to\pi^{+}\nu$. In this topology there is a single invisible particle, so one can replace the $\rho$-mass constraint by a $\tau$-mass shell constraint, $P_{\nu}+P_{\pi^{+}})^{2}\;=\;P_{\tau}^{2}\;=\;m_{\tau}^{2}$ which is analogous in spirit but not identical in implementation to the $\rho$-mass constraint used above. A variant with $\tau^{+}\to K^{+}\nu$ is also possible, but its branching fraction is substantially smaller due to CKM suppression ($\sim\!\lvert V_{us}/V_{ud}\rvert^{2}$) and reduced phase space.

### v2

@@ 1 Introduction
These displaced decays provide a unique experimental insight to search for and characterize HNLs. By reconstructing the kinematics of the decay products, searches can directly constrain the HNL mass and mixing parameters, enabling tests of both the seesaw mechanism and models of sterile neutrinos across a broad range of parameter space. Several experimental programs have placed significant constraints on the HNL parameter space across a wide mass range. In the multi-GeV regime, dedicated searches by ATLAS and CMS provide the most stringest limits on active–sterile mixing ATLAS2023 ; CMS2022 . In the sub-GeV mass range, experiments such as NA62 have recently searched for the decay $K\rightarrow\ell N$ and set upper limits on its branching fraction, thereby directly constraining HNL production in kaon decays NA62HNL2021 . A complementary search technique at running and future experiments is offered by HNL mass-peak reconstruction in stopped charged-kaon decay events Alves_2025 . Additionally, fixed-target experiments like PS191 and CHARM have been reinterpreted in modern global analyses Drewes2018 to constrain active-sterile mixing and mass range up to 2 GeV. The future experiment at a dedicated beam-dump facility at CERN, SHiP, is expected to significantly extend the search sensitivity in the mass range 0.5-5 GeV to smaller mixing angles SHIP2016 . In experimental searches, the reconstructed or inferred HNL mass is used as the final discriminant to enhance signal sensitivity and suppress background contamination. An accurate determination of the HNL mass plays a central role in setting exclusion limits and defining the sensitivity reach of each analysis across different masses and lifetimes.
@@ 5 Performance Evaluation of the $\rho$-Mass Constraint Method in Beam-Dump and Collider Scenarios
To assess the $\rho$-mass constraint method introduced in Section 4, we evaluate its performance across two representative experimental environments: (i) a fixed-target beam-dump scenario, exemplified by the SHiP experiment, and (ii) a high-energy proton-proton collider scenario, such as the LHC. These two classes of experiments differ significantly in terms of HNL production modes, event topologies, and detector geometries. The objective is twofold: first, to verify whether the kinematic constraint method is able to reconstruct a mass peak as closely as possible to the truth HNL mass and also of smaller width than the simple invariant mass $m_{\ell\pi}$; and second, to test its discriminatory power in differentiating these from $N\to\ell\pi$ that yields the same charged particles final state but lacks the additional $\pi^{0}$. This section presents simulation-based studies for each experimental scenario at particle level, hence without the effects from detector response.
@@ 5.1 Fixed-Target Scenario
Figure 2: Longitudinal neutral–pion momentum $p_{\pi^{0}}^{\parallel}$. The filled (red) histogram shows the nominal (particle level), while the hatched (blue/green) histograms show the two kinematic solutions $q_{\min}^{\parallel}$ and $q_{\max}^{\parallel}$ obtained with the $\rho$-mass constraint method. The sample corresponds to $N\!\to\!\mu\rho$ with $m_{N,GeN}=1~\mathrm{GeV}$ and $c\tau=50~\mathrm{m}$, generated with EventCalc--SHiP in a SHiP-like configuration. The $q_{\min}^{\parallel}$ solution is softer and peaks at low momentum, whereas $q_{\max}^{\parallel}$ is broader and extends to higher momenta.
@@ 5.2 Collider Scenario
Figure 4: Longitudinal neutral pion momentum $p_{\pi^{0}}^{\parallel}$ for an LHC-like sample with $m_{N,Gen}=2~\mathrm{GeV}$, produced via $pp\!\to\!W^{\pm}\!+\!X\to\ell N\!+\!X$ and $N\!\to\!\ell\rho$. The filled histogram shows the nominal (particle level), while the hatched histograms correspond to the two kinematic solutions $q_{min/max}^{\parallel}$ obtained with the $\rho$-mass constraint method. The $q_{\min}^{\parallel}$ solution is softer and peaks at low momentum, whereas $q_{\max}^{\parallel}$ is broader and extends to higher momenta.
@@
Figure 6 compares the $m_{N}^{\rho}$ distributions for the $\mu\rho$ and $\mu\pi$ HNL deccay channels. The reconstructed $m_{N}^{\rho}$ forms a displaced, narrower peak for $\mu\pi$ HNL deccay, enabling clear separation of the decay hypotheses. The observed peak separation reflects the method’s sensitivity to the underlying decay structure. At detector level, however, the achievable discrimination will be limited by the track–momentum resolution at the displaced vertex. For example, ATLAS reports percent-level dimuon mass resolution in the $\mathrm{GeV}$ range for prompt, combined muons ATLASMuonCalib2023 ; resolutions for displaced tracks are typically worse and analysis-dependent. Hence, while the particle-level peak offset is plausibly resolvable, a definitive assessment requires full detector simulation. n addition to the $m_{N}^{\rho}$ peak position and the impact parameters at the primary interaction vertex (similar to the SHiP case; see subsection 5.1), there are two other handles to distinguish the $\ell\rho$ and $\ell\pi$ HNL decay channels (in tracking-only reconstruction):
@@ 6 $W$-Mass Constraint Method
As it was mentioned in Section 3 to further reconstruct the kinematics of the HNL decay, it is useful to exploit momentum conservation in an appropriately chosen reference frame, where we follow essentially the method used for $pp\to\ell W,W\to\ell N,N\to\ell\ell^{\prime}\nu$ described in ATLAS2023 , by replacing the $\nu$ by the $\pi^{0}$ in the $N\to\ell\rho,\rho\to\pi\pi^{0}$. In the rotated frame the energy-momentum conservation relation can be written as:
@@
Analogous to the $\rho$-mass constraint, the $W$-mass constraint yields two kinematic solutions for the HNL mass, denoted as $m_{N}^{\text{W,min/max}}$ . The performance of both solutions is studied using MC-simulated events in Sec. 6.2. In some events, however, no real solution exists because the radicand in Equation 9 becomes negative. To address this, we use an adaptive $W$-mass treatment described in Section 6.1, in analogy to Ref. ATLAS2023 for $pp\to\ell W,W\to\ell N,N\to\ell\ell^{\prime}\nu$.
@@ 7 Outlook: Mass Reconstruction of HNLs Coupling to $\tau$ Leptons
Probing HNL couplings to $\tau$ leptons in HNL decays to final states with a tau lepton plus additional particles is compelling yet challenging. Unlike the electron or muon cases, the $\tau$ must be reconstructed through its decays, which reduces sensitivity due to the $\tau$ branching fractions to the used channels. Moreover, for almost all subsequent $\tau$-lepton decays, leptonic or semileptonic, the reconstruction of a mass peak is not possible due to the missing momentum from the emitted neutrino(s). A notable extension of the method applies for $m_{N}>m_{\tau}+m_{\pi}$, namely the channel $N\to\tau^{+}\pi^{-}$ with $\tau^{+}\to\pi^{+}\nu$. In this topology there is a single invisible particle, so one can replace the $\rho$-mass constraint by a $\tau$-mass shell constraint, $(P_{\nu}+P_{\pi^{+}})^{2}\;=\;P_{\tau}^{2}\;=\;m_{\tau}^{2}$ which is analogous in spirit but not identical in implementation to the $\rho$-mass constraint used above. A variant with $\tau^{+}\to K^{+}\nu$ is also possible, but its branching fraction is substantially smaller due to CKM suppression ($\sim\!\lvert V_{us}/V_{ud}\rvert^{2}$) and reduced phase space.